OpenStudy (anonymous):

How do you get 3x^2+7y^2+24x-70y=-202 into the form (x+4)^2/7+(y-5)^2/3=1?

myininaya (myininaya):

\[3x^2+24x+7y^2-70y=-202\] \[3(x^2+8x)+7(y^2-10y)=-202\] \[3(x^2+8x+(\frac{8}{2})^2)+7(y^2-10x+(\frac{10}{2})^2)=-202+3(\frac{8}{2})^2+7(\frac{10}{2})^2\] \[3(x^2+8x+4^2)+7(y^2-10x+5^2)=-202+3(16)+7(25)\] \[3(x+4)^2+7(y-5)^2=-202+48+175\] \[3(x+4)^2+7(y-5)^2=21\] \[\frac{3}{21}(x+4)^2+\frac{7}{21}(y-5)^2=\frac{21}{21}\] \[\frac{1}{7}(x+4)^2+\frac{1}{3}(y-5)^2=1\]

myininaya (myininaya):

completing the square dave is all you really need to do this trick

OpenStudy (anonymous):

Ah, I see. I totally missed that. Thanks!

myininaya (myininaya):

np